# Generalized Label-Efficient 3D Scene Parsing via Hierarchical Feature Aligned Pre-Training and Region-Aware Fine-tuning

### v1

@@ 4.4 Data-Efficient 3D Instance Segmentation
As our method can be integrated seamlessly into various network backbones and applied to different highly-level understanding tasks, we have also integrated our method with Point-Group [90] for the instance segmentation on ScanNet with results shown in Table V. Notice that the performance increase is 21.7% when merely 1% data is labeled compared with the sup-only case. It further demonstrates that our proposed approaches for the unsupervised branch have effectively exploited the unlabeled data to improve the feature learning capacity of the model. As shown in Fig. 5, our proposed WS3D and WS3D++ both provide explicit boundary guidance for separating diverse kinds of semantic classes, and the instance segmentation performance with very limited labeling percentage is comparable to those fully supervised counterparts.
@@ 4.5 Data-Efficient 3D Object Detection
For the data-efficient 3D object detection, following our previous work [99], we evaluate current approaches extensively on SUN RGB-D [137] and ScanNet [20] benchmarks for 3D object detection tasks with the strong 3D object detection backbone VoteNet [126]. It can be demonstrated the open-vocabulary designs can to some extent boost the 3D object detection performance, which demonstrate the generalization capacity of our fundation model. As shown in Table III, we have also tested the ablated approach WS3D-Open which abandons our feature-aligned pre-training term $\mathcal{L}^{total}_{Ctr}$ and directly uses the CLIP [9] feature encoder for the knowledge distillation loss term with $\mathcal{L}^{Dist}_{\mathcal{KL}}$. It can be demonstrated that the performance degradation can obviously be observed when comparing WS3D-Open to WS3D, which demonstrate the effectiveness of our proposed hierarchical feature aligned pre-training in improve data efficiency.
TABLE VI: WS3D and WS3D++ ablation studies on ScanNet (Left Value), S3DIS (Middle Value) and Nuscene (Right Value) validation set, for semantic segmentation (Metric: mIoU%) and on ScanNet validation set for the instance segmentation (Metric: AP@50%), both tested with the 5% labeled case.
Cases Base $w_{i,j}$ $H_{i,j}$ in EF $H_{i,j}$ in UCSL UCSL MS-UCSL SCE WS3D mIoUWS3D++ mIoUWS3D AP@50% WS3D++ AP@50% WS3D mAPWS3D++ mAP%
No. 1 56.2 / 54.6 / 53.9 62.8 / 58.1 / 56.3 45.6 54.8 45.8 58.9
No. 2 51.0 / 49.3 / 47.8 57.1 / 52.5 / 50.9 39.9 48.5 40.6 53.6
No. 3 49.9 / 47.2 / 46.6 53.1 / 50.9 / 47.9 37.0 44.5 38.1 51.7
No. 4 51.6 / 52.1 / 47.9 56.1 / 54.8 / 47.7 40.1 46.7 41.6 51.5
No. 5 51.1 / 51.4 / 49.6 57.1 / 53.8 / 49.8 40.7 47.6 40.8 53.6
No. 6 52.5 / 50.9 / 50.3 56.7 / 53.9 / 49.5 42.2 48.7 39.8 53.8
No. 7 49.3 / 48.0/ 47.6 55.6 / 53.5/ 50.8 38.1 46.8 39.8 54.3
No. 8 54.3 / 52.8 / 50.8 60.6 / 55.3 / 52.7 42.9 48.1 41.2 52.7
No. 9 48.1 / 45.3 / 45.7 50.8 / 47.6 / 50.6 34.8 39.8 33.9 45.3
TABLE VII: Ablation studies of the designed models of our proposed WS3D++. It is demonstrated that the proposed proposal generation approach and local and global contrastive approach can significantly boost final recognition performance.
Cases 2D Proposal 3D Proposal $\mathcal{L}^{G}_{Ctr}$ $\mathcal{L}^{L}_{Ctr}$ mIoUAP@50% ScanNet B10/N9 Nuscene B6/N9
No. 1 62.7 / 58.1 / 56.3 58.5 / 55.6 / 53.7 65.2 / 78.9 / 58.8 69.7 / 71.7 / 66.8
No. 2 55.2 / 51.9 / 50.6 51.3 / 48.8 / 47.9 56.9 / 70.7 / 53.8 66.1 / 66.6 / 64.5
No. 3 53.6 / 50.6 / 49.6 52.6 / 48.5 / 47.6 52.6 / 48.5 / 47.6 53.6 / 48.5 / 47.6
No. 5 53.2 / 51.7 / 49.3 50.2 / 46.1 / 45.8 50.8 / 45.8 / 45.5 50.8 / 46.7 / 45.8
No. 6 51.2 / 49.2 / 49.1 49.2 / 45.8 / 45.2 48.8 / 44.7 / 44.2 48.2 / 43.8 / 43.2
4.6 Ablation Study
Ablations: The results for WS3D are summarized in Table VI. We have ablated network modules in all combinations of settings as follows. Take the ScanNet instance segmentation at AP@50% as examples: Case 1: The full WS3D. Case 2: Removing the boundary prediction network, and not using the guidance of $w_{i,j}$. The framework still consists of the supervised branch, unsupervised guidance of the energy function based on the predicted confident pseudo label, and contrastive learning. This setting leads to a significant drop of 5.7% on AP. Case 3: Removing the pairwise term in the energy-based optimization function $E_{sum}$, the AP drops largely by 8.6%. Case 4: Removing $H_{i,j}$ in the energy function, the performance drops by 5.5%. Case 5: Removing $H_{i,j}$ in the unsupervised contrastive learning, the performance drops by 4.9%. Case 6: Conducting contrastive learning only with the region-level feature $\textbf{f}_{g,5}$ and $\textbf{f}_{h,5}$ at the fifth network stage, rather than at multiple stages. The performance drops by 3.4%. Case 7: Removing the unsupervised region-level contrastive learning branch, the performance drops largely by 7.5%. Case 8: Removing the supervised learning branch with the cross-entropy loss, the performance drops by 2.7%.
The results for WS3D++ are summarized in Table VII. Case 9: Only using the supervised branch, the ins. seg. performance drops significantly by 10.8%. Case 10: As demonstrated in Table VII, dropping the 2D or 3D proposals will both result in the performance drops both for the semantic segmentation as well as instance segmentation. Case 11: As shown in Table VII, either removing the global contrastive loss or the local contrastive loss results in less effective contrastive representations and affects the final downstream scene understanding performance.
Analyses: From the above ablations, some important findings are summarized: For WS3D, firstly, not using our designed modules results in a significant performance drop (Cases No. 3, No. 7, and No. 9), which demonstrates the effectiveness of the proposed unsupervised branch and learning strategies to leverage the unlabeled data. Secondly, our proposed learning strategies with boundary label $w_{i,j}$ (Case No. 2), energy function design (No. 3), high-confidence prediction based energy function design (No. 4), high-confidence based region-level contrastive learning strategy (No. 5), multi-stage contrastive learning network design (No. 6), all have a boost on the overall semantic/instance segmentation performance. The results demonstrate that the proposed energy loss is significant for semantic/instance seg. performances, because semantic boundary labels are crucial for identifying diverse objects. Thirdly, removing the supervision (Case No. 8), our method still maintains performance with a slight drop of performance by 2.7%. It further validates the robustness and feature learning capacity of our approach. Last but not least, as demonstrated in Table VII, for our proposed WS3D++, either dropping the 2D/3D proposals, or dropping the global/local contrastive loss result in considerable performance drops. It validates the effectiveness of our designs in enhancing discriminative 3D representations.
4.7 Qualitative and Quantitative Results of the Open-world Recognition Approaches
@@
In this Subsection, we further evaluate the performance of the open-world recognition capacity of our proposed approach. The results of open-world recognition are shown in Table IV. We have compared our work with the previous approach PLA [85] in establishing the sufficient point-language associations for the open-world robot learning. The results demonstrate that our proposed approach has superior performance in open-world recognition. We directly use the settings in the PLA [85] and split the categories on ScanNet [20] and Nuscene [35] into base and novel categories. It can also be validated that WS3D-Open, which abandons our feature-aligned pre-training and directly use the CLIP [9] feature encoder, provides slightly inferior performance compared with WS3D++, validating the effectiveness of our language-3D matching strategy designs. WS3D++ exhibits superior performance in terms of the open-vocabulary few-shot learning for diverse partitioning of original and novel classes. The open-world recognition results are shown in Figure 9 and Figure 11. It can be demonstrated that better foreground object awareness can be effectively capture by our proposed WS3D++ compared with PLA [85], with superior segmentation performance guided by the textual prompts. The superior open-world recognition performance can be achieved while conducting open-world learning in diverse spliting of based and novel classes, including B15/N4, B12/N7, B10/N9 for the ScanNet [20] as well as B12/N3 and B10/N5 for the NuScenes [35]. It demonstrates robustness of our proposed approach. Also, as demonstrated in Figure 10, the WS3D++ language driven-3D scene segmentation results are very precise as shown qualitatively, which is corresponding to the object queried by the language, and it demonstrates that the inference can be done based on the object, material, properties, affordance, room type, etc. It demonstrates that our proposed WS3D++ can enable the scene-level object recognition based on the semantic language queries. As further shown in Figure 12, through our effective rendering techniques, which establish the explicit 2D-3D association, the aligned representation of 2D-3D-language co-embeddings can be learned and the objectness can also be enhanced through finding the similarity among diverse views through contrastive learning approaches. Also, by combining 2D and 3D region proposals, more complete and apparent object-level information can be clearly captured both from 2D views and 3D views.
@@ 4.8 Instance Discrimination Capacity
We show t-SNE visualizations of the learned latent feature representations for various semantic classes in Fig. 13. The case study task is the semantic segmentation on the S3DIS dataset with a supervision level of 5%. It is demonstrated that more distinctive and better separated point-wise feature embeddings are provided by our proposed unsupervised region-level contrastive learning, which can be attributed to its strong instance discrimination capacity. And more separated feature space can be provided with our proposed WS3D++ compared to WS3D and GPC. This strong instance discrimination capacity can be explained by more discriminative feature representations guided by 3D vision-language aligned representations, and is thus more beneficial to high-level semantic and instance segmentation performances both in terms of data efficiency and open-world recognition capacity. Also, our proposed hierarichical feature alignment also provides more separated feature space, which means that the feature alignment successfully enhances the final instance discrimination capacity.
@@ 5 Conclusion
In this paper, we proposed a general WS3D++ framework for data-efficient 3D scene parsing. It involves both the pre-training and the fine-tuning stages. In the pre-training stage, we propose the hierarchical feature alignment strategy to acquire accurate regional 3D-linguistic pairs, thus the performance can be enhanced to a large extent. At the same time, we propose an unsupervised boundary-aware energy-based loss and a novel region-level multi-stage semantic contrastive learning strategy, which are complementary to each other to make the network learn more meaningful and discriminative features from the unlabeled data. The effectiveness of our approach is verified across three diverse large-scale 3D scene understanding benchmarks under various experiment circumstances. Our approach can maximally exploit the unlabeled data to enhance the performance both for 3D point clouds semantic segmentation and instance segmentation, under various labeling percentages in the limited reconstruction case. Our proposed label-efficient learning framework, termed WS3D++, provides conprehensive baselines for future 3D scene parsing methods when the label is inaccessible or limited.
Robot Arm Grasping Example
We deploy our approach for the task of open-world perception in robot grasping. We use our proposed approach for segmentation and use the ROS 2 Gazebo-based framework to implement the other components of the system, such as kinematics/dynamics modeling, motion planning, low-level control, point cloud-based pose estimation, etc. Our proposed approach has robust performance and decent accuracy in grasping, which demonstrates the potential of our proposed WS3D++ in industrial manipulation applications in grasping and dropping novel objects beyond the training set. The video demos are attached.
Supplementary Material
In this supplementary material, additional experimental results and details that are not included in the main paper due to space limits are provided as follows:
  • Details of our experimental datasets (see Section A).
  • Additional qualitative experimental results (see Section B).
  • The PFH-based point cloud over-segmentation method to provide the initial over-segmentation results for our proposed WS3D (see Section C).
  • Details of our data augmentation approach (see Section D).
  • Training details of the boundary prediction network (see Section E).
  • The low density issues for the outdoor point clouds captured by LiDAR sensors (see Section F).
A Experimental Datasets Details
To demonstrate the effectiveness of our proposed WS3D and WS3D++, we tested it on both indoor and outdoor 3D scene segmentation benchmarks, including S3DIS[21], SemanticKITTI[18], and ScanNet[20] for semantic segmentation, and ScanNet[20] for instance segmentation. The detailed information of each dataset is listed as follows:
Figure 14: Qualitative semantic segmentation results on SemanticKITTI with 5% labels, compared with the fully-supervised arts Cylinder3D [119], and BAAF-Net [89] with semantics indicated by diverse colors. The red circles highlight the performance difference between diverse methods.
ScanNet-V2 is a standard large-scale and widely-acknowledged 3D indoor scene understanding benchmark consisting of more than 1,600 3D scene scans. It contains densely reconstructed point clouds from RGB-D images captured by the depth camera. For the dataset partition in fully supervised learning, we follow the official partition of ScanNet-V2[20] using 1,201 scans as the training set, 312 scans as the validation set, and 100 scans as the test set.
SemanticKITTI[18] is a large-scale point cloud understanding benchmark for self-driving applications. The point cloud of individual scenes is obtained densely by Velodyne VLP-64 LiDAR. The dataset covers long-range road scenes, thanks to the vehicle driving in complex road scenarios with mounted LiDAR to capture the point clouds. More than 43,550 labeled LiDAR scans in total are split into 21 sequences. Each LiDAR scan contains approximately $10^{5}$ points. We follow the official setting[18] that uses the sequences 00 to 07, and 09 to 10 for training, the sequence 08 for validation, and the sequence 11 to 21 for testing.
S3DIS is a commonly used indoor 3D scene understanding benchmark that contains various indoor scenes. It includes six indoor areas which are made up of 271 rooms. Each room contains a magnitude of $10^{6}$ points. The typical room size is $20\times 15\times 5$ meters. In our experiments, we follow GPC[121] using area 5 as the validation set and the other areas as the training set.
NuScene is a commonly used autonomous driving dataset containing more than 1000 scenes. Every scene is elaborately annotated with the 3D bounding boxes for 23 categories and 8 attributes. For the base/novel category partition, we overlook the ’otherflat’ class and randomly divide the rest 15 classes into B12/N3 and B10/N5 as in PLA [85].
Figure 15: Results of our proposed PFH-based over-segmentation in Algorithm 1. Diverse over-segmented regions are indicated by diverse colors.
B Additional Experimental Results
We provide the additional experimental qualitative results on ScanNet and SemanticKITTI semantic segmentation and ScanNet instance segmentation.
In Fig. 14, the results on SemanticKITTI[18] are shown for the 5% labels case compared with the fully-supervised SOTAs Cylinder3D[119], and BAAF-Net[89]. From the qualitative results, our method can sometimes offer better segmentation performance than fully-supervised SOTAs for objects such as roads, trees, etc., which demonstrates the effectiveness of our proposed WS3D for outdoor 3D semantic segmentation.
Comparisons with SOTAs in the fully supervised mode. To demonstrate the feature learning capacity of our proposed WS3D, we have also experimented with the 100% labeling percentage for a fair comparison with fully supervised SOTAs. The results are shown in Table VIII and the last column of Table I. We have fed the whole training set into the supervised branch and unsupervised branch in Fig. 3 simultaneously. Therefore, our proposed energy-based loss and region-level contrastive learning strategy operate as additional optimization guidance for network training. Table VIII demonstrates that we can realize at least comparable or even better results, compared with the fully supervised SOTAs.
Transductive learning. Similar to the experimental setting of [121], we have also conducted experiments evaluating the performance of WS3D in transductive learning. Different from the inductive learning we tested above that requires the trained model to be generalized to an unseen test set, transductive learning can exploit the testing set when training. Compared with inductive learning, we add the test set as part of the unlabeled data in transductive learning. As is demonstrated in Table III, the sem. seg. mIOU becomes higher if the network is learned in a transductive way, both for the fully labeled case with 100% labels and the weakly labeled case with 20% labels. It demonstrates that our proposed WSL approaches, including the energy-based boundary-aware loss and the region-level contrastive learning, can leverage the unlabeled data for feature learning effectively in an implicit way to enhance the final segmentation performance.
Figure 16: In the limited reconstruction case, our proposed approaches WS3D and WS3D++ work well for the low-density SemanticKITTI over-segmentation and boundary prediction (5% label percentage). It can also be demonstrated that the explicit boundary information can be clearly captured, and our proposed approach is robust to the density variation of indoor dense and outdoor sparse point clouds.
C Point Clouds Over-segmentation
C.1 PFH-based Point Clouds Over-segmentation
As mentioned in Subsection 3.1 in the main paper, the previous over-segmentation methods depended on the region growing[34], which relied heavily on the accurate normal estimation and were easily influenced by noises. In our work, we use the normal, curvature, and point feature histogram[116] simultaneously to provide the initial over-segmentation results for obtaining different regions. The point feature histogram (PFH)[116] can be regarded as a robust representation of local geometric features, which is robust to 6D pose transformation and deals well with diverse sampling densities, and neighboring noises. We use it as a more robust geometric feature to obtain the over-segmentation result.
To be more specific, we denote the average normal vector of a point as $\textbf{n}=(n_{x},n_{y},n_{z})$, the PFH feature as $\textbf{f}_{pfh}\in\mathbb{R}^{125\times 1}$. The cosine of the angle between the average normal vectors of two adjacent points $P_{i}$ and $P_{j}$ is denoted as $A_{n}$, and the cosine of the angle between two PFH vectors of two adjacent points in the 125-dimensional space is denoted as $A_{pfh}$. The Affinity $A(P_{i},P_{j})$ of two adjacent points $P_{i}$ and $P_{j}$ is given as:
$A(P_{i},P_{j})=\sqrt{A_{n}^{2}+A_{pfh}^{2}}.$ (14)
According to the Point Cloud Library (PCL)[34], in the process of region growing, points with an affinity larger than a threshold $T_{thres}$ are merged iteratively. We set $T_{thres}$ as 1 following the typical setting[34]. As shown in Algorithm 1, we summarize the detailed procedures for the PFH-based over-segmentation as the following steps:
Input: The input raw point set $\textbf{P}_{i}=\{p_{i}\},i=1,2,...,N_{i}$. $p_{i}=(x_{i},y_{i},z_{i})$.
Output: The output pseudo label matrix $\textbf{{L}}_{region}$ for different regions.
Initialize $P_{seed}=p_{\{seed,i\}},i=1,2,...,N^{seed}$;
Initialize the pseudo label matrix $\textbf{{L}}_{region}$ as a zero matrix. while not converged do
       Select $K$ nearest neighbour points $p_{j}$ around the seed points $p_{seed}$ for comparisons based on fast Octree-based $K$ Nearest Neighbour search;
       for the seed points $p_{j}$ selected do
             if Condition 1 then
                  Assign $p_{j}$ the same overseg-label as $p_{seed}$;
                   if Condition 2 then
                        Regard the point $p_{j}$ as new seed points;
            else
                  Assign $p_{j}$ with a new overseg-label. Regard $p_{j}$ as new seed points;
            $j\leftarrow j+1$;
             Update the pseudo label matrix $\textit{{L}}_{region}$.
return The class label matrix $\textit{{L}}_{region}$ of $\textbf{P}_{i}$ with different clusters.
Algorithm 1 The Proposed PFH-based Region Growing Over-segmentation to Obtain Regions
  1. 1.
    Select the points that have top 2‰ minimum curvatures[34] as the initial seed points $p_{seed}$. Utilizing fast $K$ Nearest Neighbour search[120], we obtain the $j_{th}$ neighbouring point $p_{j}$ of the initial seed point $p_{seed}$, and calculate the Affinity $A(p_{seed},p_{j})$ between $p_{seed}$ and $p_{j}$.
  2. 2.
    If the Affinity $A(P_{seed},P_{j})$ is larger than $T_{thres}$ (Condition 1), we assign the $j_{th}$ neighbouring point $p_{j}$ with the same label as $p_{seed}$.
  3. 3.
    If Condition 1 is satisfied, denote the curvature of $p_{j}$ and $p_{seed}$ as $r_{j}$ and $r_{seed}$, and denote the difference of $r_{j}$ and $r_{seed}$ as $\Delta r=\|r_{j}-r_{seed}\|$. If $\Delta r\leq\sigma$ (Condition 2), we set $p_{j}$ as the seed point. That means we set the points $p_{j}$ as seed points, only if the difference in curvature is small enough. Otherwise, we only assign $p_{j}$ the same label as $p_{seed}$.
  4. 4.
    If Condition 1 is not satisfied, we assign the neighbouring point $p_{j}$ with a different label from $p_{seed}$, and also assign $p_{j}$ as the seed point.
Finally, points are assigned to initial over-segmented regions with labels. The simple algorithm is summarized in Algorithm 1. The loop will terminate if any one of the following convergence conditions is satisfied, which are proposed:
  1. 1.
    All points have been assigned with labels;
  2. 2.
    There are no seed points that can be added.
After the procedures above, we can obtain the label $\textit{{L}}_{region}$, which indicates the separated regions obtained by over-segmentation. Some randomly selected over-segmentation results are shown in Fig. 15. The PFH-based over-segmentation can automatically divide the whole point cloud scene into geometrically separated regions as shown in Fig. 15. Different colors indicate different regions.
TABLE VIII: Comparison of SOTAs methods in the semantic segmentation performance on ScanNet validation set, S3DIS validation set (Area 5), and on SemanticKITTI validation set and test set. All results are based on the 100% label ratio. Top-two results are highlighted.
Approaches Venue ScanNet S3DIS SemanticKITTI
Val. Val. (Area 5) Val. Test
Minkow-Network [37] CVPR19 72.2 65.4 61.1 63.1
PointASNL [49] CVPR20 66.4 62.6 - 58.8
KPConv [144] ICCV19 69.2 67.1 - 46.8
SPV-NAS [145] ECCV20 - - 64.7 66.4
Fusion-Net ECCV20 - 67.2 63.7 61.3
MV-Fusion [44] ECCV20 76.4 65.4 - -
Cylinder3D [119] CVPR21 - - 65.9 67.8
BAAF-Net [89] CVPR21 - 72.1 61.2 59.9
Sup-only-GPC [121] ICCV21 70.9 66.4 65.0 65.4
GPC [121] ICCV21 74.0 68.8 65.8 67.7
WS3D ECCV22 76.9 72.9 66.9 69.0
WS3D++ - 77.6 73.5 67.7 69.8
D Data Augmentation Details
D.1 Data Augmentation Approaches
We perform three kinds of data augmentations to enhance the unlabeled data and to perform region-level contrastive learning, including: 1. Geometric rotations and mirroring of the point cloud scene. 2. Adding minor disturbance to the point cloud scene. 3. Random scaling of the point cloud scene.
[…]
Adding minor disturbance. We add the noise for the normalized point clouds coordinates, which can be represented as $P_{x}^{d}=P_{x}^{m}+T_{d}$. $T_{d}$ is the Gaussian noise with the mean of 0 and variance of 0.01. We set $T_{d}$ in the range of $[-0.1,0.1]$.
Random scaling. We do random scaling to the point cloud scene. Random scaling of ($0.9-1.1$) is done to the point clouds coordinates. Denote the scaling parameter as $\eta$, the final augmented point clouds scene $P_{x}^{{}^{\prime}}$ is given as $P_{x}^{{}^{\prime}}=\eta P_{x}^{d}$.
E Training Details of the Boundary Prediction Network
[…]
F Low-Density Issues of Outdoor LiDAR points
While using the boundary prediction network to extract the semantic boundary for large-scale 3D scenes, the density issue should be taken into consideration because the density variation may influence the quality of the obtained pseudo labels. Randomly selected results of SKITTI LiDAR over-segmentation and boundary predictions results are shown in Fig. 16. The density influences over-segmentation performance especially for outdoor low-density point clouds because the 3D point cloud normals, curvatures, and PFH features are all directly related to density. For outdoor sparse SemanticKITTI, we have a very simple but effective trick to tackle low-density outdoor LiDAR points: discarding all regions containing less than $N_{thres}$ points, i.e. not using these regions in all our network modules. ($N_{thres}=100$ was set empirically for SKITTI). As shown in Fig. 16, isolated regions with fewer points than $N_{thres}$ are successfully discarded for reliable and robust region-level predictions. As shown in Fig. 16, compared to Karpathy’s approach, our PFH-based over-segmentation better distinguishes similar semantic classes such as road and sidewalk, revealing that our adapted PFH can have a better local feature description capacity than other classical features, both for our proposed WS3D and WS3D++. The results of boundary predictions are not much or greatly influenced by density as shown in Fig. 16. Our boundary prediction network learns high-level semantic boundaries, which are not explicitly related to the low-level point density. In summary, our proposed WS3D and WS3D++ still work well for the low-density cases.

### v2

@@ 4.4 Data-Efficient 3D Instance Segmentation
As our method can be integrated seamlessly into various network backbones and applied to different highly-level understanding tasks, we have also integrated our method with Point-Group [73] for the instance segmentation on ScanNet with results shown in Table VI. Notice that the performance increase is 21.7% when merely 1% data is labeled compared with the sup-only case. It further demonstrates that our proposed approaches for the unsupervised branch have effectively exploited the unlabeled data to improve the feature learning capacity of the model. Our proposed WS3D and WS3D++ both provide explicit boundary guidance for separating diverse kinds of semantic classes, and the instance segmentation performance with very limited labeling percentage is comparable to those fully supervised counterparts.
@@ 4.5 Data-Efficient 3D Object Detection
For the data-efficient 3D object detection, following our previous work [81], we extensively evaluate current approaches extensively on SUN RGB-D [111] and ScanNet [18] benchmarks for 3D object detection tasks with the strong 3D object detection backbone VoteNet [115]. It can be demonstrated the open-vocabulary designs can to some extent boost the 3D object detection performance, which demonstrate the generalization capacity of our fundation model. As shown in Table V, we have also tested the ablated approach WS3D-Open which abandons our feature-aligned pre-training term $\mathcal{L}^{total}_{Ctr}$ and directly uses the CLIP [7] feature encoder for the knowledge distillation loss term with $\mathcal{L}^{Dist}_{\mathcal{KL}}$. It can be demonstrated that the performance degradation can be observed when comparing WS3D-Open with WS3D++, which demonstrate the effectiveness of our proposed hierarchical feature aligned pre-training in improving the data efficiency in downstream scene parsing.
4.6 Qualitative and Quantitative Results of the Open-world 3D Recognition Approaches
@@
In this Subsection, we further evaluate the performance of the open-world recognition capacity of our proposed approach. The results of open-world recognition are shown in Table I. We have also compared our work with the previous approach PLA [69] in establishing the sufficient point-language associations for the open-world robot learning. The results demonstrate that our proposed approach has superior performance in open-world recognition. We directly use the settings in the PLA [69] and split the categories on ScanNet [18] and Nuscene [36] into base and novel categories. It can also be validated that WS3D-Open, which abandons our feature-aligned pre-training and directly use the CLIP [7] feature encoder, provides slightly inferior performance compared with WS3D++, validating the effectiveness of our language-3D matching strategy designs. WS3D++ exhibits superior performance in terms of the open-vocabulary few-shot learning for diverse partitioning of original and novel classes. The open-world recognition results are shown in Figure 9 and Figure 10. It can be demonstrated that better foreground object awareness can be effectively capture by our proposed WS3D++ compared with PLA [69], with superior segmentation performance guided by the textual prompts. The superior open-world recognition performance can be achieved while conducting open-world learning in diverse spliting of based and novel classes, including B15/N4, B12/N7, B10/N9 for the ScanNet [18] as well as B12/N3 and B10/N5 for the NuScenes [36]. It demonstrates robustness of our proposed approach. Also, as demonstrated in Figure 11, the WS3D++ language driven-3D scene segmentation results are very precise as shown qualitatively, which is corresponding to the object queried by the language, and it demonstrates that the inference can be done based on the object, material, properties, affordance, room type, etc. It demonstrates that our proposed WS3D++ can enable the scene-level object recognition based on the semantic language queries. As further shown in Figure 13, through our effective rendering techniques, which establish the explicit 2D-3D association, the aligned representation of 2D-3D-language co-embeddings can be learned and the object information can also be enhanced through finding the similarity among diverse views through contrastive learning approaches. Also, by combining 2D and 3D region proposals, more complete and apparent object-level information can be clearly captured both from 2D views and 3D views. It turns out that in the first place, while initializing other various weakly supervised approaches, our proposed approach can realize consistent improvement on the final performance of weakly supervised scene parsing, which demonstrates the superior generalization capacity benefiting from our designed pre-training of our proposed WS3D++ framework. In the second pace, the performance of our proposed WS3D++ is comparatively superior compared with the existing weakly supervised comparative approaches listed above.
@@ 4.7 Instance Discrimination Capacity
We show t-SNE visualizations of the learned latent feature representations for various semantic classes in Fig. 12. The case study task is the semantic segmentation on the S3DIS dataset with a supervision level of 5%. It is demonstrated that more distinctive and better separated point-wise feature embeddings are provided by our proposed unsupervised region-level contrastive learning, which can be attributed to its strong instance discrimination capacity. And more separated feature space can be provided and maintained with our proposed WS3D++ compared to WS3D and GPC. This strong instance discrimination capacity can be explained by more discriminative feature representations guided by 3D vision-language aligned representations, and is thus more beneficial to high-level semantic and instance segmentation performances both in terms of data efficiency and open-world recognition capacity. Also, our proposed hierarichical feature alignment also provides more separated feature space, which means that the feature alignment successfully enhances the final instance discrimination capacity.
Robot Arm Grasping Example. We have deployed our approach for the task of open-world perception in robot grasping in our extended robotic research work. We use our proposed approach for segmentation and use the ROS 2 Gazebo-based framework to implement the other components of the system, such as kinematics/dynamics modeling, motion planning, low-level control, point cloud-based pose estimation, etc. Our proposed approach has robust performance and decent accuracy in grasping, which demonstrates the potential of our proposed WS3D++ in industrial manipulation applications in grasping and dropping novel objects beyond the training set.
@@ 5 Conclusion
In this paper, we proposed a general WS3D++ framework for open-vocabulary and data-efficient 3D scene parsing. The whole framework involves both the pre-training and the fine-tuning stages. During the pre-training stage, we propose the hierarchical feature alignment strategy to acquire accurate regional 3D-linguistic pairs, thus the performance can be enhanced to a large extent. At the same time, we propose an unsupervised boundary-aware energy-based loss and a novel region-level multi-stage semantic contrastive learning strategy, which are complementary to each other to make the network learn more meaningful and discriminative features from the unlabeled data. The effectiveness of our approach is verified across three diverse large-scale 3D scene understanding benchmarks under various experiment circumstances. Our approach can maximally exploit the unlabeled data to enhance the performance both for 3D point clouds semantic segmentation and instance segmentation, and object detection under various labeling percentages in the limited reconstruction case. Our proposed label-efficient learning framework, termed WS3D++, provides conprehensive baselines for future 3D scene parsing methods when the label is inaccessible or limited. The proposed pre-training as well as fine-tuning approach can have a significant boost on the final open-vocabulary and data-efficient semantic scene parsing in term of efficiency, effectiveness, and robustness.